### v1

# On Polyconvexity and Almgren Uniform Ellipticity With Respect to Polyhedral Test Pairs

Abstract
We study anisotropic geometric energy functionals defined on a class of $k$-dimensional surfaces in a Euclidean space. The classical notion of ellipticity, coming from Almgren, for such functionals is investigated. We extend a recent result of De Rosa, Lei, and Young and show that uniform ellipticity of an anisotropic energy functional with respect to real polyhedral chains implies uniform polyconvexity of the integrand.
1 Introduction
Let $\mathbf{G}_{0}(n,k)$ denote the Grassmannian of oriented $k$-planes in $\mathbf{R}^{n}$. Given a continuous integrand $F:\mathbf{G}_{0}(n,k)\to(0,\infty)$, one defines anisotropic energy functional $\Phi_{F}:\mathbf{P}_{k}(\mathbf{R}^{n})\to\mathbf{R}$, acting on $k$-dimensional polyhedral chains (see §4 for relevant definitions) in $\mathbf{R}^{n}$ by
$\Phi_{F}(T)=\int F(\vec{T}(x))\,\mathrm{d}\|T\|(x)\quad\text{for }T\in\mathbf{P}_{k}(\mathbf{R}^{n}),$
where $\|T\|$ is a Radon measure over $\mathbf{R}^{n}$ and $\vec{T}$ is a $\|T\|$-measurable function with values in $\mathbf{G}_{0}(n,k)$.
Assuming $F$ is even (i.e. $F(-\xi)=F(\xi)$ for $\xi\in\mathbf{G}_{0}(n,k)$) we may consider $F$ to be defined on the unoriented Grassmannian and then the functional $\Phi_{F}$ can be extended to act on varifolds by setting $\Phi_{F}(V)=\int F(T)\,\mathrm{d}V(x,T)$ for a $k$-dimensional varifold $V$ in $\mathbf{R}^{n}$; see [All72]. Almgren studied the regularity of varifolds which are local minimisers of $\Phi_{F}$ and proved that they are almost everywhere regular, provided $F$ is uniformly elliptic (abbreviated AUE) [Alm68, Theorems 1.4 and 1.7]. If $F$ is elliptic, then a flat $k$-disc (or $k$-cube) is the unique minimiser of $\Phi_{F}$ among all competitors with the same ($k-1$)-dimensional boundary as the disc (or cube). Uniform ellipticity gives additionally quantitative control on $\Phi_{F}$; see §5.
This notion heavily depends on the choice of competitors. When competitors are rectifiable currents, any uniformly convex norm on ${\textstyle\bigwedge}_{k}\mathbf{R}^{n}$ gives rise to a uniformly elliptic integrand; cf. [Fed69, 5.1.2]. Burago and Ivanov [BI04, Theorem 3] showed that if competitors are rectifiable currents (with real coefficients), then semi-ellipticity is equivalent to being extendibly convex, i.e., there exists an extension of the integrand to a norm on the whole ${\textstyle\bigwedge}_{k}\mathbf{R}^{n}$. Their proof relies on showing, in non-constructive way, the existence of a polyhedral surface with prescribed tangent planes; see [BI04, Theorem 1, Theorem 2].
However, if competitors are allowed only to be integral currents, then there exists a continuous integrand defined on $\mathbf{G}_{0}(4,2)$ for which the associated energy functional is elliptic, but the integrand cannot be extended to a norm on the whole ${\textstyle\bigwedge}_{2}\mathbf{R}^{4}$; see [BI04, Theorem 5].
In [DLY24, Theorem 3.2], De Rosa, Lei, and Young provided a constructive proof of [BI04, Theorem 2] and established an analogous theorem to [BI04, Theorem 3] for Lipschitz multivalued functions, namely, that ellipticity is equivalent to polyconvexity; see [DLY24, Theorem 1.7].
The classical notion of polyconvexity, coming from calculus of variations, is defined for functionals acting on functions; cf. [Giu03, Chapter 5]. In this setting integrands are real-valued functions defined on the space of linear maps $\operatorname{Hom}(\mathbf{R}^{k},\mathbf{R}^{n})$ – for the purposes of the following discussion let us call them classical integrands as opposed to geometric integrands of the type $\mathbf{G}_{0}(n,k)\to\mathbf{R}$; cf. [Fed69, 5.1.1 and 5.1.9].
The authors of [DLY24] use a variant of polyconvexity for classical integrands and briefly describe its connection with being extendibly convex for related geometric integrands. Later they (as we do as well) call a geometric integrand $F$ polyconvex in case it is extendibly convex. As a service to the community, we present a detailed formal discussion of the relation between extendibly convex geometric integrands and polyconvex classical integrands in section 3.
A notion stronger than polyconvexity is uniform polyconvexity (abbreviated UPC) which involves certain constant $c>0$; see 3.21. In this article, we show that uniform ellipticity for polyhedral $k$-chains with constant $c$ implies uniform polyconvexity with any constant smaller than $c$. This can be seen as a refinement of some results contained in [BI04] and [DLY24].
Theorem (see 5.10 for a precise statement).
Assume $F$ is a geometric integrand satisfying $\operatorname{Lip}F<\infty$. If $F$ is Almgren uniformly elliptic with respect to polyhedral $k$-chains with real coefficients and with ellipticity constant $0<c<\infty$, then $F$ is uniformly polyconvex with any constant $0<\tilde{c}<c$.
Our proof is based on [DLY24, Theorem 3.2], which allows to construct polyhedral $k$-chains with Gaussian measures close to a given measure and provides control of the geometry of the constructed surface. Note also that the reverse implication (with $\tilde{c}=c$) holds by a simple argument involving the Stokes theorem; cf. [Fed69, 5.1.2].
This paper is organised as follows. In Section 2, we recall the basic terminology that will be used throughout the paper. In Section 3, we present a formal discussion on the relationship between the notions of polyconvex and extendibly convex integrands, demonstrating their equivalence. Moreover, in the same section, we recall the definition of uniform polyconvexity for integrands. Section 4 is devoted to the classical definitions of polyhedral chains and Gaussian measures. Finally, in Section 5, we define ellipticity for polyhedral test pairs and prove the main result.
2 Preliminaries
In principle we shall follow the notation of Federer; see [Fed69, pp. 669–671]. In particular, $A\!\sim\!B$ is the set-theoretic difference and $\mathscr{P}$ denotes the set of positive integers. Diverting from Federer, if $-\infty<s<r<\infty$, we occasionally write $(s,r)$, $[s,r]$, $(s,r]$, and $[s,r)$ to denote appropriate intervals in $\mathbf{R}$. We shall use letters $k$ and $n$ to denote two integers satisfying $0\leq k<n$. We say that a function $f$ defined on some open subset of a Banach space with values in another Banach space is of class $\mathscr{C}^{l}$ if it is $l$-times differentiable and $\mathrm{D}^{l}f$ is continuous; cf [Fed69, 3.1.11].
2.1 Definition (cf. [Fed69, 1.3.2]).
By $\Lambda(n,k)$ we denote set of all increasing maps of $\{1,\ldots,k\}$ into $\{1,\ldots,n\}$.
Remark.
Note that any $\lambda\in\Lambda(n,k)$ is uniquely determined by $\operatorname{im}\lambda$ which is a $k$-element subset of $\{1,2,\ldots,n\}$; thus, choosing $\lambda\in\Lambda(n,k)$ corresponds to choosing $k$-elements out of $\{1,2,\ldots,n\}$.
2.2 Definition.
Let $\xi\in{\textstyle\bigwedge}_{k}\mathbf{R}^{n}$. We define the vector space associated with $\xi$ by
$\operatorname{space}\xi=\mathbf{R}^{n}\cap\bigl{\{}v:v\wedge\xi=0\bigr{\}}\,.$
2.3 Remark ([Fed69, 1.6.1]).
An element $\xi\in{\textstyle\bigwedge}_{k}\mathbf{R}^{n}$ is simple if and only if $\dim\operatorname{space}\xi=k$.
2.4 Definition ([Fed69, 3.2.28(b)]).
We define the oriented Grassmannian $\mathbf{G}_{0}(n,k)$ as
$\mathbf{G}_{0}(n,k)={\textstyle\bigwedge}_{k}\mathbf{R}^{n}\cap\bigl{\{}\xi:\xi\text{ is simple},\,|\xi|=1\bigr{\}}\,.$
2.5 Remark.
Let $\xi\in{\textstyle\bigwedge}_{k}\mathbf{R}^{n}$. Recall from [Fed69, 1.8.1] that $\|\xi\|=|\xi|$ if and only if $\xi$ is simple.
2.6 Definition.
Given a vectorspace $X$ we write $\mathrm{End}({X})$ for the space of all linear maps mapping $X$ to $X$, i.e., $\mathrm{End}({X})=\operatorname{Hom}(X,X)$.
2.7 Definition ([Fed69, 1.7.4]).
A linear map $j\in\operatorname{Hom}(\mathbf{R}^{k},\mathbf{R}^{n})$ is called an orthogonal injection if $j^{*}\circ j=\mathds{1}_{\mathbf{R}^{k}}$. If $p\in\operatorname{Hom}(\mathbf{R}^{n},\mathbf{R}^{k})$ and $j=p^{*}$ is an orthogonal injection, then we say that $p$ is an orthogonal projection. The set of all orthogonal projections $\mathbf{R}^{n}\to\mathbf{R}^{k}$ is denoted $\mathbf{O}^{\ast}({n},{k})$.
2.8 Remark.
To each $\lambda\in\Lambda(n,k)$ corresponds the map $\operatorname{\textbf{p}}_{\lambda}\in\mathbf{O}^{\ast}({n},{k})$ defined as
$\operatorname{\textbf{p}}_{\lambda}(x)=(x_{\lambda(1)},\ldots,x_{\lambda(k)})\quad\text{for $x=(x_{1},\ldots,x_{n})\in\mathbf{R}^{n}$}\,.$
2.9 Definition (cf. [Roc70, p. 14]).
Let $X$ be a vectorspace and $A\subseteq X$. We define
$\operatorname{ray}A=X\cap\{ta:0\leq t<\infty,\,a\in A\}\quad\text{and}\quad\operatorname{cone}A=\operatorname{conv}(\operatorname{ray}A)\,.$
2.10 Definition.
Let $X$ be a normed vectorspace and $f:X\to\mathbf{R}$. We say that $f$ is positively homogeneous if $f(tx)=tf(x)$ whenever $x\in X$ and $0\leq t<\infty$.
2.11 Definition.
Let $X$ be a vectorspace and $f:X\to\mathbf{R}$. We say that $f$ is a gauge if it is convex non-negative and positively homogeneous. We say that $f$ is a strict gauge if $f$ is a gauge and
$f(x+y)<f(x)+f(y)\quad\text{for all $x,y\in X\!\sim\!\{0\}$ such that $x\notin\operatorname{ray}\{y\}$}\,.$
2.12 Definition.
A signed measure over $\mathbf{G}_{0}(n,k)$ is a Daniell integral on the lattice of functions $\mathscr{K}(\mathbf{G}_{0}(n,k))$ as defined in [Fed69, 2.5.6].
2.13 Definition ([Fed69, 2.5.5]).
Total variation of a signed measure $\mu$ over $\mathbf{G}_{0}(n,k)$ is defined as
$\|\mu\|_{\mathrm{TV}}:=|\mu|(\mathbf{G}_{0}(n,k)),\quad\text{where}\quad|\mu|=\mu^{+}+\mu^{-}\,.$
3 Polyconvexity
3.1.
The notion of polyconvexity has been introduced, in the context of calculus of variations, for integrands defined on the space of linear maps $\operatorname{Hom}(\mathbf{R}^{k},\mathbf{R}^{n})$; see, e.g., [Giu03, Ch.5, p. 147]. In our case integrands are functions on $\mathbf{G}_{0}(n,k)\subseteq{\textstyle\bigwedge}_{k}\mathbf{R}^{n}$ and the corresponding notion of polyconvexity can be found, e.g., in [DLY24, 1.3]. The relation between these two notions has been very briefly explained in a small remark after Definition 1.3 in [DLY24]. As a service to the community, as well as for the sake of author’s own understanding, we present here a detailed formal discussion of this relation.
3.2 Remark.
Assume $f\in\operatorname{Hom}(\mathbf{R}^{k},\mathbf{R}^{n})$ and $A\in\mathbf{R}^{n\times k}$ is the matrix of $f$ written in the standard bases. For any $\lambda\in\Lambda(n,k)$ the matrix of the map $\operatorname{\textbf{p}}_{\lambda}\circ f$ is obtained by choosing the rows of $A$ indexed by $\lambda$ and $\det(\operatorname{\textbf{p}}_{\lambda}\circ f)$ is a $k\times k$ minor of the matrix $A$. Having this in mind one readily sees that the following two definitions are a special case of the ones given by Giusti.
3.3 Definition.
Let $f\in\operatorname{Hom}(\mathbf{R}^{k},\mathbf{R}^{n})$. We define $M(f)\in\mathbf{R}^{\Lambda(n,k)}$ by
$M(f)(\lambda)=\det(\operatorname{\textbf{p}}_{\lambda}\circ f)\quad\text{for $\lambda\in\Lambda(n,k)$}\,.$
3.4 Definition (cf. [Giu03, Ch.5, p. 147]).
A function $G:\operatorname{Hom}(\mathbf{R}^{k},\mathbf{R}^{n})\to[0,\infty)$ is called (strictly) polyconvex if there exists a (strict) gauge $h:\mathbf{R}^{\Lambda(n,k)}\to[0,\infty)$ such that $G=h\circ M$.
3.5 Remark.
Note that our definition differs from [Giu03, Ch.5] in two ways. First, we assume $G$ depends only on minors of top degree $k$ and not on all minors. Second, we enforce $h$ to be positively homogeneous, which immediately yields that $G$ is positively $k$-homogeneous and determines the growth of $G$ at infinity.
3.6 Definition.
An integrand is a continuous non-negative function $F:\mathbf{G}_{0}(n,k)\to\mathbf{R}$.
3.7 Remark.
Note that an integrand $F$ is defined only on the Grassmannian $\mathbf{G}_{0}(n,k)$. Later it will be useful to consider positively homogeneous extensions of $F$ onto the whole of ${\textstyle\bigwedge}_{k}\mathbf{R}^{n}$.
3.8 Remark.
Let $(e_{1},\ldots,e_{k})$ be the standard basis of $\mathbf{R}^{k}$, $E=e_{1}\wedge\cdots\wedge e_{k}\in\mathbf{G}_{0}(k,k)$, $\xi\in\mathbf{G}_{0}(n,k)$, $p\in\mathbf{O}^{\ast}({n},{k})$ be such that ${\textstyle\bigwedge}_{k}p^{*}E=\xi$, and $A\in\mathbf{R}^{n\times k}$ be the matrix of $p^{*}$ in the standard bases. Then the columns of $A$ form an orthonormal basis of $\operatorname{space}\xi$ which is positively oriented with respect to $\xi$. This explains the connection of the following definition with the one given by De Rosa, Lei, and Young.
3.9 Definition (cf. [DLY24, 1.3]).
Let $(e_{1},\ldots,e_{k})$ be the standard basis of $\mathbf{R}^{k}$. An integrand $F$ is said to be (strictly) polyconvex if there is a (strict) gauge
$h:\mathbf{R}^{\Lambda(n,k)}\to[0,+\infty)\quad\text{such that}\quad F(\xi)=h(M(p_{\xi}^{*}))$
whenever $\xi\in\mathbf{G}_{0}(n,k)$ and $p_{\xi}\in\mathbf{O}^{\ast}({n},{k})$ satisfies ${\textstyle\bigwedge}_{k}p_{\xi}^{*}(e_{1}\wedge\cdots\wedge e_{k})=\xi$.
3.10 Remark.
Definition 3.9 is independent of the choice of $p_{\xi}$ for $\xi\in\mathbf{G}_{0}(n,k)$. Indeed, let $\xi\in\mathbf{G}_{0}(n,k)$ and $p_{1},p_{2}\in\mathbf{O}^{\ast}({n},{k})$ be such that ${\textstyle\bigwedge}_{k}p_{i}^{*}(e_{1}\wedge\cdots\wedge e_{k})=\xi$ for $i\in\{1,2\}$. There exists $T\in\mathbf{O}({k})$ such that $p_{1}^{*}=p_{2}^{*}\circ T$ and we get
$M(p_{1}^{*})(\lambda)=\det(\operatorname{\textbf{p}}_{\lambda}\circ p_{1}^{*})=\det(\operatorname{\textbf{p}}_{\lambda}\circ p_{2}^{*}\circ T)=\det(\operatorname{\textbf{p}}_{\lambda}\circ p_{2}^{*})\det T=M(p_{2}^{*})(\lambda)\det T$
for $\lambda\in\Lambda(n,k)$. However, we also know that
$\xi={\textstyle\bigwedge}_{k}p_{2}^{*}(e_{1}\wedge\cdots\wedge e_{k})={\textstyle\bigwedge}_{k}p_{1}^{*}(e_{1}\wedge\cdots\wedge e_{k})={\textstyle\bigwedge}_{k}(p_{2}^{*}\circ T)(e_{1}\wedge\cdots\wedge e_{k})\\ ={\textstyle\bigwedge}_{k}p_{2}^{*}\circ{\textstyle\bigwedge}_{k}T(e_{1}\wedge\cdots\wedge e_{k})=\det T{\textstyle\bigwedge}_{k}p_{2}^{*}(e_{1}\wedge\cdots\wedge e_{k})\,;$ (1)
hence, $\det T=1$.
3.11 Remark.
Let $e_{1},\ldots,e_{k}\in\mathbf{R}^{k}$ be the standard basis of $\mathbf{R}^{k}$, $E=e_{1}\wedge\cdots\wedge e_{k}\in\mathbf{G}_{0}(k,k)$, $\xi\in\mathbf{G}_{0}(n,k)$, and $p\in\mathbf{O}^{\ast}({n},{k})$ be such that ${\textstyle\bigwedge}_{k}p^{*}E=\xi$. In 3.10 we showed that $M(p)$ does not depend on the particular choice of $p$ so $M$ could actually be defined on $\mathbf{G}_{0}(n,k)$. Let $u_{1},\ldots,u_{n}$ be the standard basis of $\mathbf{R}^{n}$. For $\lambda\in\Lambda(n,k)$ set $u_{\lambda}=u_{\lambda(1)}\wedge\cdots\wedge u_{\lambda(k)}$ and note that $u_{\lambda}={\textstyle\bigwedge}_{k}\operatorname{\textbf{p}}_{\lambda}^{*}E$; thus,
$\xi\bullet u_{\lambda}={\textstyle\bigwedge}_{k}p^{*}E\bullet{\textstyle\bigwedge}_{k}\operatorname{\textbf{p}}_{\lambda}^{*}E={\textstyle\bigwedge}_{k}(p_{\lambda}\circ p^{*})E\bullet E=\det(p_{\lambda}\circ p^{*})=M(p^{*})(\lambda)\,.$
3.12 Definition.
Define $N:{\textstyle\bigwedge}_{k}\mathbf{R}^{n}\to\mathbf{R}^{\Lambda(n,k)}$ by the formula
$N(\xi)(\lambda)=\xi\bullet u_{\lambda}\quad\text{for $\xi\in{\textstyle\bigwedge}_{k}\mathbf{R}^{n}$ and $\lambda\in\Lambda(n,k)$}\,.$
3.13 Remark.
Note that $N$ is a linear isometry.
3.14 Corollary.
Let $F$ be an integrand. The following are equivalent
  1. (a)
    $F$ is (strictly) polyconvex in the sense of definition 3.9.
  2. (b)
    There exists of (strict) gauge $h:\mathbf{R}^{\Lambda(n,k)}\to[0,\infty)$ such that $F(\xi)=h(N(\xi))$ for $\xi\in\mathbf{G}_{0}(n,k)$.
3.15 Corollary.
Assume $h:\mathbf{R}^{\Lambda(n,k)}\to[0,\infty)$. The following are equivalent
  1. (a)
    $h\circ M:\operatorname{Hom}(\mathbf{R}^{k},\mathbf{R}^{n})\to\mathbf{R}$ is (strictly) polyconvex in the sense of 3.4;
  2. (b)
    $h\circ N:{\textstyle\bigwedge}_{k}\mathbf{R}^{n}\to\mathbf{R}$ is (strictly) polyconvex in the sense of 3.9;
  3. (c)
    $h|N[\mathbf{G}_{0}(n,k)]$ can be extended to a (strict) gauge on $\mathbf{R}^{\Lambda(n,k)}$.
3.16 Remark.
We could have defined polyconvexity in 3.4 and 3.9 without enforcing $h$ to be positively homogeneous (only convex) and we would get an analogous corollary to 3.15. Observe, however, that if $h:\mathbf{R}^{\Lambda(n,k)}\to[0,\infty)$ is only convex, then, in general, it might not be possible to find a gauge which agrees with $h$ on $N[\mathbf{G}_{0}(n,k)]$; hence, not assuming $h$ to be positively homogeneous in 3.9 significantly affects the class of polyconvex integrands.
To support this claim let $\mathbb{S}=\mathbf{R}^{\Lambda(n,k)}\cap\{y:|y|=1\}$, choose $e\in N[\mathbf{G}_{0}(n,k)]\subseteq\mathbb{S}$, $1<r<\infty$, and $(1+r)^{2}/r<p<\infty$ (note that $(1+r)^{2}/r\geq 4$). Consider the non-negative convex function $h:\mathbf{R}^{\Lambda(n,k)}\to[0,\infty)$ given by
$h(x)=|x-re|^{p}\quad\text{for $x\in\mathbf{R}^{\Lambda(n,k)}$}\,.$
Next, define a non-negative positively homogeneous function $f:\mathbf{R}^{\Lambda(n,k)}\to[0,\infty]$ by setting
$f(x)=h(x)\quad\text{and}\quad f(tx)=tf(x)\quad\text{for $x\in\mathbb{S}$ and $0\leq t<\infty$}\,.$
Fix $x=-e$ and $u\in\mathbb{S}$ such that $e\perp u$. Choose a geodesic $\gamma:\mathbf{R}\to\mathbb{S}$ such that $\gamma(0)=x$, $\gamma^{\prime}(0)=u$ (then, necessarily $\gamma^{\prime\prime}(0)=-x$). Direct computations show
$(f\circ\gamma)^{\prime\prime}(0)=\mathrm{D}^{2}f(x)uu-\mathrm{D}f(x)x=\mathrm{D}^{2}f(x)uu-f(x)=\mathrm{D}^{2}f(x)uu-h(x)\,,$ (2)
$(f\circ\gamma)^{\prime\prime}(0)=(h\circ\gamma)^{\prime\prime}(0)=\mathrm{D}^{2}h(x)uu-\mathrm{D}h(x)x\,,$ (3)
$\mathrm{D}h(x)x=p|x-re|^{p-2}(x-re)\bullet x\,,$ (4)
$\text{and}\quad\mathrm{D}^{2}h(x)uu=p(p-2)|x-re|^{p-4}\bigl{(}(x-re)\bullet u\bigr{)}^{2}+p|x-re|^{p-2}\,;$ (5)
hence, recalling $x=-e$ and $e\perp u$,
$\mathrm{D}^{2}f(x)uu=\mathrm{D}^{2}h(x)uu-\mathrm{D}h(x)x+h(x)=|x-re|^{p-2}\bigl{(}p-p(x-re)\bullet x+|x-re|^{2}\bigr{)}\\ =|x-re|^{p-2}\bigl{(}-pr+(1+r)^{2}\bigr{)}\,.$ (6)
Since $p>(1+r)^{2}/r$ we get $\mathrm{D}^{2}f(x)uu<0$ so $f$ is not convex.
Convex hulls
3.17 Definition (cf. [Fed69, 2.5.19]).
Assume $X$ is a set and $x\in X$. The Dirac measure over $X$ having a single atom at $x$ is denoted by
$\boldsymbol{\delta}_{x}(A)=\begin{cases}1\text{\quad if \quad}x\in A\\ 0\text{\quad if \quad}x\notin A\end{cases}\quad\text{for $A\subseteq X$}\,.$
3.18 Definition.
Let $F$ be an integrand. The convex positively homogeneous hull of $F$ is defined by
$\operatorname{Conv}_{F}(\xi)=\inf\left\{\int{}{}F\,\mathrm{d}\mu:\begin{gathered}N\in\mathscr{P},\,m_{1},\ldots,m_{N}\in(0,\infty),\,\eta_{1},\ldots,\eta_{n}\in\mathbf{G}_{0}(n,k),\,\\ \mu={\textstyle\sum_{i=1}^{N}}m_{i}\boldsymbol{\delta}_{\eta_{i}},\,\quad\xi={\textstyle\int}\eta\,\mathrm{d}\mu(\eta)\end{gathered}\right\}\\ \text{for $\xi\in{\textstyle\bigwedge}_{k}\mathbf{R}^{n}$}\,.$ (7)
3.19 Remark.
$\operatorname{Conv}_{F}$ is the biggest convex positively homogeneous function (i.e. a gauge) not exceeding $F$ on $\mathbf{G}_{0}(n,k)$ pointwise.
3.20 Lemma.
An integrand $F$ is polyconvex if and only if $F=\operatorname{Conv}_{F}|_{\mathbf{G}_{0}(n,k)}$.
Proof.
Assume $F$ is polyconvex so that there exists a gauge $h:\mathbf{R}^{\Lambda(n,k)}\to\mathbf{R}$ such that $F=h\circ N|\mathbf{G}_{0}(n,k)$. Since $N$ is a linear isometry, we see that $h\circ N$ is a gauge on ${\textstyle\bigwedge}_{k}\mathbf{R}^{n}$ so
$F=h\circ N|\mathbf{G}_{0}(n,k)\leq\operatorname{Conv}_{F}|\mathbf{G}_{0}(n,k)\leq F\,.$
On the other hand, if $\operatorname{Conv}_{F}|\mathbf{G}_{0}(n,k)=F$, we set $h=\operatorname{Conv}_{F}\circ N^{-1}$ to get $F=h\circ N|\mathbf{G}_{0}(n,k)$. ∎
Uniform polyconvexity
3.21 Definition.
Let $0<c<\infty$ and $F$ be an integrand. We say that $F$ is uniformly polyconvex (with constant $c$) if
${\textstyle\sum_{i=1}^{d}}m_{i}F(\eta_{i})-F(\eta_{0})\geq c\bigl{(}{\textstyle\sum_{i=1}^{d}}|\eta_{i}|-|\eta_{0}|\bigr{)}$
whenever $d\in\mathscr{P}$, $m_{1},\ldots,m_{d}\in(0,\infty)$, $\eta_{0},\ldots,\eta_{d}\in\mathbf{G}_{0}(n,k)$ are such that $\eta_{0}={\textstyle\sum_{i=1}^{d}}m_{i}\eta_{i}$.
If $F$ is uniformly polyconvex with constant $c$, we shall write $F\in\mathrm{UPC(c)}$.
3.22 Remark.
Assume $F\in\mathrm{UPC(c)}$ for some $0<c<\infty$, $d\in\mathscr{P}$, $m_{1},\ldots,m_{d}\in(0,\infty)$, $\eta_{0},\ldots,\eta_{d}\in\mathbf{G}_{0}(n,k)$, and $\eta_{0}={\textstyle\sum_{i=1}^{d}}m_{i}\eta_{i}$. Set $\mu={\textstyle\sum_{i=1}^{N}}m_{i}\boldsymbol{\delta}_{\eta_{i}}$. Then we have
${\textstyle\int}F\,\mathrm{d}\mu={\textstyle\sum_{i=1}^{d}}m_{i}F(\eta_{i})\geq F(\eta_{0})\,;$
hence, $\operatorname{Conv}_{F}(\eta_{0})=F(\eta_{0})$. Since $\eta_{0}$ was arbitrary, we see from 3.20 that $F$ is polyconvex. Consequently, we get that uniformly polyconvexity implies polyconvexity.
4 Polyhedral Chains
4.1.
For the convenience of the reader and the sake of completeness we recall here the classical definitions concerning polyhedral chains.
4.2 Definition ([Fed69, 4.1.7]).
Let $U$ be an open subset of $\mathbf{R}^{n}$ and $m$ be nonnegative integer. By $\mathscr{D}^{m}(U)$ we denote set of differential forms of degree $m$ and class $\mathscr{C}^{\infty}$ on $U$ with real coefficients, whose support is a compact subset of $U$. By $\mathscr{D}_{m}(U)$ we denote set of $m$-dimensional currents in $U$, i.e., the space of continuous linear functionals on $\mathscr{D}^{m}(U)$.
4.3 Remark ([Fed69, 4.1.5 and 4.1.7]).
In case $T\in\mathscr{D}_{m}(U)$ is representable by integration, then $T=\|T\|\wedge\vec{T}$, where $\|T\|$ is a Radon measure over $U$ and $\vec{T}$ is a $\|T\|$-measurable function with values in ${\textstyle\bigwedge}_{m}\mathbf{R}^{n}$.
4.4 Definition ([Fed69, 4.1.8]).
Let $A$ and $B$ be open subsets of $\mathbf{R}^{m}$ and $\mathbf{R}^{n}$. Moreover, let
$p:A\times B\to A\quad\text{and}\quad q:A\times B\to B$
be the canonical projections onto the first and second factor respectively. For any $S\in\mathscr{D}_{i}(A)$ and $T\in\mathscr{D}_{j}(B)$ the Cartesian product of $S$ and $T$ is the current
$S\times T\in\mathscr{D}_{i+j}(A\times B)$
characterised by the following conditions
$(S\times T)(p^{\#}\alpha\wedge q^{\#}\beta)=\left\{\begin{aligned} &S(\alpha)T(\beta)&&\text{if $k=i$}\\ &0&&\text{if $k\neq i$}\end{aligned}\right.\quad\text{for $\alpha\in\mathscr{D}^{k}(A)$ and $\beta\in\mathscr{D}^{i+j-k}(B)$}\,.$
4.5 Definition ([Fed69, 4.1.11]).
Let $H:\mathbf{R}^{n}\times\mathbf{R}\times\mathbf{R}^{n}\to\mathbf{R}^{n}$ be a map such that
$H(x,t,y)=(1-t)x+ty\quad\text{whenever}\quad x,y\in\mathbf{R}^{n}\quad\text{and}\quad t\in\mathbf{R}\,.$
For any currents $S\in\mathscr{D}_{i}(\mathbf{R}^{n})$ and $T\in\mathscr{D}_{j}(\mathbf{R}^{n})$ with compact supports, the join of $S$ and $T$ is defined as
$S\mathchoice{\times\mskip-15.5mu\times}{\times\mskip-15.5mu\times}{\times\mskip-15.5mu\times}{\times\mskip-15.5mu\times}T=H_{\#}(S\times[0,1]\times T)\in\mathscr{D}_{i+1+j}(\mathbf{R}^{n})$
4.6 Definition ([Fed69, 4.1.11]).
Let $u_{0},\ldots,u_{k}\in\mathbf{R}^{n}$. An oriented $k$-simplex $\llbracket u_{0},\ldots,u_{m}\rrbracket$ is defined inductively as
$\llbracket u\rrbracket=\boldsymbol{\delta}_{u}\text{ whenever $u\in\mathbf{R}^{n}$}\,,$ (8)
$\llbracket u_{0},\ldots,u_{k}\rrbracket=\llbracket u_{0}\rrbracket\mathchoice{\times\mskip-15.5mu\times}{\times\mskip-15.5mu\times}{\times\mskip-15.5mu\times}{\times\mskip-15.5mu\times}\llbracket u_{1},\ldots,u_{k}\rrbracket\in\mathscr{D}_{k}(\mathbf{R}^{n})\quad\text{if $k\in\mathscr{P}$}\,.$ (9)
If $u_{0},\ldots,u_{k}$ are affinely independent, we say that $\llbracket u_{1},\ldots,u_{k}\rrbracket$ is non-degenerate.
4.7 Remark ([Fed69, 4.1.11]).
Every oriented $k$-simplex in any Euclidean space is representable as affine image of any non-degenerate oriented $k$-simplex in $\mathbf{R}^{k}$.
4.8 Remark.
Let $\sigma\subseteq\mathbf{R}^{n}$ be the convex hull of $u_{0},\ldots,u_{k}\in\mathbf{R}^{n}$ and $\tau:\sigma\to\mathbf{G}_{0}(n,k)$ be the orientation of $\sigma$ defined by
$\tau(x)=\frac{(u_{1}-u_{0})\wedge\ldots\wedge(u_{k}-u_{k-1})}{|(u_{1}-u_{0})\wedge\ldots\wedge(u_{k}-u_{k-1})|}\quad\text{for $x\in\sigma$}\,.$
Then
$\llbracket u_{0},\ldots,u_{k}\rrbracket(\omega)={\textstyle\int_{\sigma}}\langle\omega(x),\tau(x)\rangle\,\mathrm{d}\mathscr{H}^{k}(x)\quad\text{for $\omega\in\mathscr{D}^{k}(U)$}\,.$
4.9 Definition ([Fed69, 4.1.22]).
Let $U$ be an open subset of $\mathbf{R}^{n}$, and $K\subseteq U$ be compact. The additive subgroup of $\mathscr{D}_{k}(U)$ generated by all oriented $k$-simplices $\llbracket u_{0},\ldots,u_{k}\rrbracket$ such that $\operatorname{conv}\{u_{0},\ldots,u_{k}\}\subseteq K$ is denoted
$\mathscr{P}_{k,K}(U)\,.$
All the $k$-dimensional integral polyhedral chains in $U$ form the abelian group
$\mathscr{P}_{k}(U)={{\textstyle\bigcup}}\bigl{\{}\mathscr{P}_{k,K}(U):K\subseteq U\text{ compact}\bigr{\}}\,.$
4.10 Definition ([Fed69, 4.1.22]).
Let $\mathbf{P}_{k,K}(U)$ be the vector-subspace of $\mathscr{D}_{k}(U)$ generated by $\mathscr{P}_{k,K}(U)$. The $k$-dimensional polyhedral chains in $U$ are members of the vectorspace
$\mathbf{P}_{k}(U)={{\textstyle\bigcup}}\bigl{\{}\mathbf{P}_{k,K}(U):K\subseteq U\text{ compact}\bigr{\}}\,.$
4.11 Remark.
Elements of $\mathbf{P}_{k}(U)$ are precisely finite sums of the form
[…]
Furthermore, we may assume $a_{i}>0$ for $i\in\{1,2,\ldots,m\}$ since if $a_{i}<0$ for some $i\in\{1,2,\ldots,m\}$, then we obtain the same $T$ by replacing $a_{i}$ with $-a_{i}$ and reversing the orientation of $\Delta_{i}$. In particular, any polyhedral chain has compact support. Now, let $\omega\in\mathscr{D}^{k}(U)$, then
$T(\omega)={\textstyle\sum_{i=1}^{m}}a_{i}\Delta_{i}(\omega)={\textstyle\int_{\bigcup_{i=1}^{m}\operatorname{spt}\Delta_{i}}}\langle\omega(x),{\textstyle\sum_{i=1}^{m}}a_{i}\tau_{i}(x)\rangle\,\mathrm{d}\mathscr{H}^{k}(x)\,;$
thus, every polyhedral chain can be seen as rectifiable $k$-current $(E,\tau,\theta)$ corresponding to the countably $k$-rectifiable set
$E={{\textstyle\bigcup}}_{i=1}^{m}\operatorname{spt}\Delta_{i}$
with orientation given by
$\tau(x)=\frac{\tilde{\tau}(x)}{|\tilde{\tau}(x)|}\quad\text{for $x\in E$}\,,\quad\text{where}\quad\tilde{\tau}(x)={\textstyle\sum_{i=1}^{m}}a_{i}\vec{\Delta}_{i}(x)\boldsymbol{\chi}_{\operatorname{spt}\Delta_{i}}(x)$
and multiplicity function
$\theta(x)=|\tilde{\tau}(x)|\,.$
Gaussian measures
4.12 Definition ([DLY24, Ch.1]).
Let $A=(E,\tau,\theta)$ be a polyhedral $k$-chain. The weighted Gaussian image of $A$ is the measure $\gamma_{A}$ over $\mathbf{G}_{0}(n,k)$ defined by
[…]
4.13 Definition ([DLY24, Ch.1]).
Let $F$ be an integrand. For every polyhedral $k$-chain $A=(E,\tau,\theta)$ we define anisotropic energy functional $\Phi_{F}$ as
$\Phi_{F}(A)={\textstyle\int_{E}}\theta\cdot F\circ\tau\,\mathrm{d}\mathscr{H}^{k}\,.$
4.14 Remark.
By the very definition of the push-forward [Fed69, 2.1.2] the anisotropic energy functional can be written in terms of the weighted Gaussian image as follows
[…]
4.15 Definition.
The mass of $T\in\mathbf{P}_{k}(U)$ is defined as
$\mathbf{M}(T)=\|T\|(U)\,.$
4.16 Remark.
In the case when $T={\textstyle\sum_{i=1}^{m}}a_{i}\Delta_{i}\in\mathbf{P}_{k}(U)$ for some $a_{1},\ldots,a_{m}\in(0,\infty)$ and oriented $k$-simplices $\Delta_{1},\ldots,\Delta_{m}$, we have
$\mathbf{M}(T)={\textstyle\sum_{i=1}^{m}}a_{i}\mathscr{H}^{k}(\Delta_{i})=\gamma_{T}(\mathbf{G}_{0}(n,k))=\|\gamma_{T}\|_{\mathrm{TV}}\,.$
5 Ellipticity
5.1 Definition.
Recalling [Fed69, 4.1.8 and 4.1.32] for $k\in\mathscr{P}$ we set
$Q_{k}=\underbrace{\llbracket 0,1\rrbracket\times\cdots\times\llbracket 0,1\rrbracket}_{\text{$k$ factors}}\in\mathscr{P}_{k}(\mathbf{R}^{k})\,.$
5.2 Definition (cf. [RK20, 4.1]).
We say that $(S,D)\in\mathcal{P}$ is a polyhedral test pair if there exists $p\in\mathbf{O}^{\ast}({n},{k})$ such that
$S\in\mathbf{P}_{k}(U)\,,\quad D=(p^{*})_{\#}Q_{k}\,,\quad\text{and}\quad\partial S=\partial D\,.$
5.3 Definition.
Let $F$ be an integrand and $\mathcal{P}$ be the set of polyhedral test pairs. We say that $F$ is elliptic with respect to $\mathcal{P}$ if there exists $0<c<\infty$ (the ellipticity constant) such that
$\Phi_{F}(S)-\Phi_{F}(D)>c\bigl{(}\mathbf{M}(S)-\mathbf{M}(D)\bigr{)}\quad\text{for $(S,D)\in\mathcal{P}$}\,.$
If this holds we write $F\in\mathrm{AUE}(\mathcal{P},c)$.
Main result
5.4.
Our main result depends on the theorem proven recently by De Rosa, Lei, and Young. Before stating it we first discuss some definitions.
5.5 Definition (cf. [DLY24, §3]).
Let $e_{1},\ldots,e_{n}$ be the standard basis of $\mathbf{R}^{n}$ and $P\subseteq\mathbf{R}^{n}$ be a linear subspace of dimension $k$. We say that $P$ has rational slope if there exists $f\in\operatorname{Hom}(\mathbf{R}^{n},\mathbf{R}^{n-k})$ such that $P=\ker f$ and $f(e_{i})\in\mathbf{Z}^{n-k}$ for $i\in\{1,2,\ldots,n\}$.
5.6 Definition.
Let $u_{1},\ldots,u_{n-k}$ be the standard basis of $\mathbf{R}^{n-k}$ and $\ast:{\textstyle\bigwedge}_{n-k}\mathbf{R}^{n}\to{\textstyle\bigwedge}_{k}\mathbf{R}^{n}$ be the Hodge star defined with respect to the standard orientation of $\mathbf{R}^{n}$; cf. [Fed69, 1.7.8]. Set
$\xi_{f}=\ast\bigl{(}f^{*}(u_{1})\wedge\cdots\wedge f^{*}(u_{n-k})\bigr{)}\in{\textstyle\bigwedge}_{k}\mathbf{R}^{n}\quad\text{for $f\in\operatorname{Hom}(\mathbf{R}^{n},\mathbf{R}^{n-k})$}$ (10)
$\text{and}\quad Q(n,k)=\bigl{\{}t\xi_{f}:t\in\mathbf{R},\,f\in\operatorname{Hom}(\mathbf{R}^{n},\mathbf{R}^{n-k}),\,f^{*}(u_{i})\in\mathbf{Z}^{n}\text{ for }i\in\{1,2,\ldots,n-k\}\bigr{\}}\,.$ (11)
5.7 Remark.
Let $e_{1},\ldots,e_{n}$ be the standard basis of $\mathbf{R}^{n}$ and $u_{1},\ldots,u_{n-k}$ be the standard basis of $\mathbf{R}^{n-k}$. If $P\subseteq\mathbf{R}^{n}$ is a $k$-dimensional linear space with rational slope and $f\in\operatorname{Hom}(\mathbf{R}^{n},\mathbf{R}^{n-k})$ is such that $P=\ker f$ and $f(e_{i})\in\mathbf{Z}^{n-k}$ for $i\in\{1,2,\ldots,n\}$, then $\xi_{f}\in Q(n,k)$. On the other hand, if $\xi\in Q(n,k)$, then there is $f\in\operatorname{Hom}(\mathbf{R}^{n},\mathbf{R}^{n-k})$ such that $f(e_{i})\in\mathbf{Z}^{n-k}$ for $i\in\{1,2,\ldots,n\}$ and
$\ker f=(\operatorname{im}f^{*})^{\perp}=\operatorname{space}\xi\,;$
hence, $\operatorname{space}\xi$ has rational slope.
5.8 Lemma.
There holds
$S_{\mathbf{Q}}=\bigl{\{}w_{1}\wedge\cdots\wedge w_{k}:w_{1},\ldots,w_{k}\in\mathbf{Q}^{k}\bigr{\}}\subseteq Q(n,k)\subseteq S_{\mathbf{R}}={\textstyle\bigwedge}_{k}\mathbf{R}^{n}\cap\{\eta:\eta\text{ is simple}\}$
and $S_{\mathbf{Q}}$ is dense in $S_{\mathbf{R}}$.
Proof.
Let $e_{1},\ldots,e_{n}$ be the standard basis of $\mathbf{R}^{n}$ and $u_{1},\ldots,u_{n-k}$ be the standard basis of $\mathbf{R}^{n-k}$. We first show that $S_{\mathbf{Q}}$ is dense in $S_{\mathbf{R}}$. Suppose $10<\varepsilon<$ 1 and $\eta=v_{1}\wedge\cdots\wedge v_{k}$ for some $v_{1},\ldots,v_{k}\in\mathbf{R}^{n}$. Without loss of generality assume $|v_{i}|=1$ for $i\in\{1,2,\ldots,k\}$. Since $\mathbf{Q}^{n}$ is dense in $\mathbf{R}^{n}$, there exist $w_{1},\ldots,w_{k}\in\mathbf{Q}^{n}$ such that $|w_{i}-v_{i}|\leq 2^{-k}\varepsilon$ for $i\in\{1,2,\ldots,k\}$. Setting $\zeta=w_{1}\wedge\cdots\wedge w_{k}$ and $r_{i}=w_{i}-v_{i}$ for $i\in\{1,2,\ldots,k\}$ we get
$|\zeta-\eta|=|(v_{1}+r_{1})\wedge\cdots\wedge(v_{k}+r_{k})-v_{1}\wedge\cdots\wedge v_{k}|\leq{\textstyle\sum_{i=1}^{k}}{\textstyle\binom{k}{i}}\varepsilon^{i}\leq\varepsilon\,.$
Next, we show that $S_{\mathbf{Q}}\subseteq Q(n,k)$. Assume $w_{1},\ldots,w_{n}\in\mathbf{Q}^{n}$ and $(w_{1}\wedge\cdots\wedge w_{n})\bullet(e_{1}\wedge\cdots\wedge e_{n})=s\neq 0$. Set $\zeta=w_{1}\wedge\cdots\wedge w_{k}$. We shall find $t\in\mathbf{R}$ and $f\in\operatorname{Hom}(\mathbf{R}^{n},\mathbf{R}^{k})$ satisfying $f(e_{i})\in\mathbf{Z}^{n}$ for $i\in\{1,2,\ldots,k\}$ and such that $\zeta=t\xi_{f}$. Define
$y_{1}=w_{1}\,,\quad y_{j}=w_{j}-{\textstyle\sum_{i=1}^{j-1}}|y_{i}|^{-2}(w_{j}\bullet y_{i})y_{i}\quad\text{for $j\in\{2,3,\ldots,n\}$}\,.$
Note that
$|y_{i}|^{2}\in\mathbf{Q}\,,\quad y_{i}\in\mathbf{Q}^{n}\,,\quad y_{i}\perp y_{j}\,,$ (12)
$\quad\text{and}\quad y_{1}\wedge\cdots\wedge y_{i}=w_{1}\wedge\cdots\wedge w_{i}\quad\text{for $i,j\in\{1,2,\ldots,n\}$ with $i\neq j$}\,.$ (13)
Clearly there is $M\in\mathscr{P}$ such that $My_{i}\in\mathbf{Z}^{n}$ for $i\in\{1,2,\ldots,n\}$. Define $f\in\operatorname{Hom}(\mathbf{R}^{n},\mathbf{R}^{n-k})$ by requiring
$f^{*}(u_{j})=My_{k+j}\quad\text{for $j\in\{1,2,\ldots,n-k\}$}\,.$
Then, recalling [Fed69, 1.7.8 and 1.5.2] and $y_{1}\wedge\cdots\wedge y_{n}=se_{1}\wedge\cdots\wedge e_{n}$, we get
$f(e_{i})\in\mathbf{Z}^{n-k}\quad\text{for $i\in\{1,2,\ldots,n\}$}\,,$ (14)
$\xi_{f}=\ast\bigl{(}My_{k+1}\wedge\cdots\wedge My_{n}\bigr{)}=M^{n-k}(-1)^{k(n-k)}\frac{|y_{k+1}|\cdots|y_{n}|}{|y_{1}|\cdots|y_{k}|}y_{1}\wedge\cdots\wedge y_{k}=t^{-1}\zeta\,,$ (15)
where $t=M^{k-n}(-1)^{k(n-k)}|y_{1}|\cdots|y_{k}|\bigl{(}|y_{k+1}|\cdots|y_{n}|\bigr{)}^{-1}\neq 0$. ∎
5.9 Theorem (cf. [DLY24, 3.2]).
Assume
$\eta_{0}\in\mathbf{G}_{0}(n,k)\,,\quad\eta_{1},\ldots,\eta_{d}\in Q(n,k)\cap\mathbf{G}_{0}(n,k)\,,\quad m_{1},\ldots,m_{d}\in(0,\infty)\,,$ (16)
${\textstyle\sum_{i=1}^{d}}m_{i}\eta_{i}=\eta_{0}\,,\quad E\in\mathbf{G}_{0}(k,k)\text{ corresponds to the standard orientation of~{}$\mathbf{R}^{k}$}\,,$ (17)
$p\in\mathbf{O}^{\ast}({n},{k})\text{ is such that }{\textstyle\bigwedge}_{k}p^{*}(E)=\eta_{0}\,,\quad D=(p^{*})_{\#}Q_{k}\,,\quad\text{and}\quad\mu={\textstyle\sum_{i=1}^{d}}m_{i}\boldsymbol{\delta}_{\eta_{i}}\,.$ (18)
Then for any $\epsilon>0$, there exists $A\in\mathbf{P}_{k}(\mathbf{R}^{n})$ with coefficients in $\mathbf{Q}$ such that
$\partial A=\partial D\quad\text{and}\quad\|\gamma_{A}-\mu\|_{\mathrm{TV}}<\epsilon\,.$
5.10 Theorem.
Assume $F$ is a Lipschitzian integrand and $0<c_{2}<c_{1}<\infty$. If $F\in\mathrm{AUE}(\mathcal{P},c_{1})$, then $F\in\mathrm{UPC(c_{2})}$.
Proof.
We argue by contradiction. Assume $F\in\mathrm{AUE}(\mathcal{P},c_{1})$ is Lipschitzian with $\operatorname{Lip}F=L<\infty$ and $F\notin\mathrm{UPC(c_{2})}$. Use the Kirszbraun theorem [Fed69, 2.10.43] to extend $F$ to the whole of ${\textstyle\bigwedge}_{k}\mathbf{R}^{n}$ preserving the Lipschitz constant. We shall use the same symbol $F$ to denote this extension. Since $F\notin\mathrm{UPC(c_{2})}$, there exist $d\in\mathscr{P}$, $\eta_{0},\ldots,\eta_{d}\in\mathbf{G}_{0}(k,n)$, and $m_{1},\ldots,m_{d}\in(0,\infty)$ such that
${\textstyle\sum_{i=1}^{d}}m_{i}\eta_{i}=\eta_{0}\quad\text{and}\quad{\textstyle\sum_{i=1}^{d}}m_{i}F(\eta_{i})-F(\eta_{0})<c_{2}\bigl{(}{\textstyle\sum_{i=1}^{d}}m_{i}|\eta_{i}|-|\eta_{0}|\bigr{)}\,.$ (19)
Since $|\eta_{i}|=1$ for $i\in\{1,2,\ldots,d\}$, $\bigl{|}\sum_{i=1}^{d}m_{i}\eta_{i}\bigr{|}=1$, and the Euclidean norm on ${\textstyle\bigwedge}_{k}\mathbf{R}^{n}$ is strictly convex, we see that $d\geq 3$ and $\sum_{i=1}^{d}m_{i}>1$. Let
$M=\sup\operatorname{im}F\quad\text{and}\quad\epsilon=\frac{\frac{1}{2}(c_{1}-c_{2})\big{(}\sum_{i=1}^{d}m_{i}-1\big{)}}{\frac{3}{2}L+M+M\binom{n}{k}^{\frac{1}{2}}+c_{1}}\,.$
For each $i\in\{0,1,\ldots,d\}$ we choose, using 5.8, a unit simple $k$-vector $\tilde{\eta}_{i}\in Q(n,k)\cap\mathbf{G}_{0}(n,k)$ satisfying
$|\eta_{i}-\tilde{\eta_{i}}|<\frac{\epsilon}{2dm_{i}}\quad\text{if $i\geq 1$}\quad\text{and}\quad|\eta_{0}-\tilde{\eta_{0}}|<\frac{\epsilon}{2}\,;$ (20)
next, we set
$\zeta=\tilde{\eta_{0}}-{\textstyle\sum_{i=1}^{d}}m_{i}\tilde{\eta_{i}}\,.$
Since $|\eta_{i}|=1=|\tilde{\eta}_{i}|$ for $i\in\{0,1,\ldots,d\}$ we obtain
$|\zeta|=\bigl{|}\tilde{\eta_{0}}-{\textstyle\sum_{i=1}^{d}}m_{i}\tilde{\eta_{i}}\bigr{|}$ $\leq\bigl{|}\tilde{\eta_{0}}-\eta_{0}\bigr{|}+\bigl{|}\eta_{0}-{\textstyle\sum_{i=1}^{d}}m_{i}\tilde{\eta_{i}}\bigr{|}$ (21)
$=\bigl{|}\tilde{\eta_{0}}-\eta_{0}\bigr{|}+\bigl{|}{\textstyle\sum_{i=1}^{d}}m_{i}\eta_{i}-{\textstyle\sum_{i=1}^{d}}m_{i}\tilde{\eta_{i}}\bigr{|}$ (22)
$\leq\bigl{|}\tilde{\eta_{0}}-\eta_{0}\bigr{|}+{\textstyle\sum_{i=1}^{d}}m_{i}\bigl{|}\eta_{i}-\tilde{\eta_{i}}\bigr{|}$ (23)
$<\frac{\epsilon}{2}+\frac{\epsilon}{2}=\epsilon\,.$ (24)
Let $e_{1},\ldots,e_{n}$ be the standard basis of $\mathbf{Q}^{n}\subseteq\mathbf{R}^{n}$ and $e_{\lambda}=e_{\lambda(1)}\wedge\cdots\wedge e_{\lambda(k)}$ for $\lambda\in\Lambda(n,k)$. Recalling [Fed69, 1.7.5] we see that $\{e_{\lambda}:\lambda\in\Lambda(n,k)\}$ is an orthonormal basis of ${\textstyle\bigwedge}_{k}\mathbf{R}^{n}$. Set
$m_{\lambda}=|\zeta\bullet e_{\lambda}|\quad\text{and}\quad\eta_{\lambda}=\operatorname{sign}(\zeta\bullet e_{\lambda})e_{\lambda}\quad\text{for $\lambda\in\Lambda(n,k)$}\,,$ (25)
$L=\Lambda(n,k)\cap\{\lambda:m_{\lambda}>0\}\,.$ (26)
We get
$\tilde{\eta_{0}}={\textstyle\sum_{i=1}^{d}}m_{i}\tilde{\eta_{i}}+{\textstyle\sum_{\lambda\in L}}m_{\lambda}\eta_{\lambda}\,.$
Denote
$\mu={\textstyle\sum_{i=1}^{d}}m_{i}\boldsymbol{\delta}_{\eta_{i}}\quad\text{and}\quad\tilde{\mu}={\textstyle\sum_{i=1}^{d}}m_{i}\boldsymbol{\delta}_{\tilde{\eta}_{i}}+{\textstyle\sum_{\lambda\in L}}m_{\lambda}\boldsymbol{\delta}_{\eta_{\lambda}}\,.$
Clearly $\eta_{\lambda}\in Q(n,k)\cap\mathbf{G}_{0}(n,k)$ and $m_{\lambda}>0$ for $\lambda\in L$. Therefore, the set of unit simple $k$-vectors $\{\tilde{\eta}_{0}$, $\tilde{\eta}_{1}$, …, $\tilde{\eta}_{d}\}\cup\{\eta_{\lambda}:\lambda\in L\}$, the numbers $\{m_{1},\ldots,m_{d}\}\cup\{m_{\lambda}:\lambda\in L\}$, and the measure $\tilde{\mu}$ satisfy the assumptions of theorem 5.9. Thus, there exist $k$-polyhedral chains $A,\tilde{D}\in\mathbf{P}_{k}(\mathbf{R}^{n})$ such that
$\vec{D}=\tilde{\eta}_{0}\,,\quad\operatorname{spt}D\text{ is a unit cube in $\operatorname{space}\tilde{\eta}_{0}$}\,,\quad\partial A=\partial\tilde{D}\,,\quad\text{and}\quad\|\gamma_{A}-\tilde{\mu}\|_{\mathrm{TV}}<\epsilon\,.$
By the assumption that $F\in\mathrm{AUE}(\mathcal{P},c_{1})$ we have the following
$\Phi(A)-\Phi(\tilde{D})$ $={\textstyle\int_{\mathbf{G}_{0}(n,k)}}F\,\mathrm{d}\gamma_{A}-{\textstyle\int_{\mathbf{G}_{0}(n,k)}}F\,\mathrm{d}\boldsymbol{\delta}_{\tilde{\eta}_{0}}$ (27)
$={\textstyle\int_{\mathbf{G}_{0}(n,k)}}F\,\mathrm{d}(\gamma_{A}-\tilde{\mu})$ (28)
$\phantom{=}+{\textstyle\int_{\mathbf{G}_{0}(n,k)}}F\,\mathrm{d}(\tilde{\mu}-\mu)$ (29)
$\phantom{=}+{\textstyle\int_{\mathbf{G}_{0}(n,k)}}F\,\mathrm{d}\mu-F(\tilde{\eta}_{0})\,.$ (30)
We will show separately upper bounds for (28), (29) and (30).
Estimate of (28) Obviously
${\textstyle\int_{\mathbf{G}_{0}(n,k)}}F\,\mathrm{d}(\gamma_{A}-\tilde{\mu})\leq M\|\gamma_{A}-\tilde{\mu}\|_{\mathrm{TV}}<M\epsilon\,.$ (31)
Estimate of (29) Recalling (20) and (21) we get
${\textstyle\int_{\mathbf{G}_{0}(n,k)}}F\,\mathrm{d}(\tilde{\mu}-\mu)$ $={\textstyle\int_{\mathbf{G}_{0}(n,k)}}F\,\mathrm{d}\tilde{\mu}-{\textstyle\int_{\mathbf{G}_{0}(n,k)}}F\,\mathrm{d}{\mu}$ (32)
$={\textstyle\sum_{i=1}^{d}}m_{i}F(\tilde{\eta}_{i})+{\textstyle\sum_{\lambda\in L}}m_{\lambda}F(\eta_{\lambda})-{\textstyle\sum_{i=1}^{d}}m_{i}F(\eta_{i})$ (33)
$={\textstyle\sum_{i=1}^{d}}\bigl{(}m_{i}F(\tilde{\eta}_{i})-m_{i}F(\eta_{i})\bigr{)}+{\textstyle\sum_{\lambda\in L}}m_{\lambda}F(\eta_{\lambda})$ (34)
$\leq L{\textstyle\sum_{i=1}^{d}}m_{i}|\eta_{i}-\tilde{\eta}_{i}|+M{\textstyle\sum_{\lambda\in L}}m_{\lambda}$ (35)
$\leq L{\textstyle\sum_{i=1}^{d}}|m_{i}||\tilde{\eta_{i}}-\eta_{i}|+M\bigl{(}{\textstyle\sum_{\lambda\in L}}m_{\lambda}^{2}\bigr{)}^{\frac{1}{2}}(\operatorname{card}L)^{\frac{1}{2}}$ (36)
$<\tfrac{1}{2}L\epsilon+M\bigl{(}{\textstyle\sum_{\lambda\in L}}|m_{\lambda}\eta_{\lambda}|^{2}\bigr{)}^{\frac{1}{2}}{\textstyle\binom{n}{k}}^{\frac{1}{2}}$ (37)
$=\tfrac{1}{2}L\epsilon+M{\textstyle\binom{n}{k}}^{\frac{1}{2}}|\zeta|<\bigl{(}\tfrac{1}{2}L+M{\textstyle\binom{n}{k}}^{\frac{1}{2}}\bigr{)}\epsilon\,,$ (38)
where the next to last inequality is a consequence of $\{\eta_{\lambda}:\lambda\in\Lambda(n,k)\}$ being an orthonormal basis and the Pythagorean Theorem.
Estimate of (30) Recalling (19) we get
${\textstyle\int_{\mathbf{G}_{0}(n,k)}}F\,\mathrm{d}\mu-F(\tilde{\eta}_{0})$ $={\textstyle\sum_{i=1}^{d}}m_{i}F(\eta_{i})-F(\eta_{0})+(F(\eta_{0})-F(\tilde{\eta_{0}}))$ (39)
$<c_{2}\bigl{(}{\textstyle\sum_{i=1}^{d}}m_{i}|\eta_{i}|-|\eta_{0}|\bigr{)}+L|\eta_{0}-\tilde{\eta_{0}}|$ (40)
$<c_{2}\bigl{(}{\textstyle\sum_{i=1}^{d}}m_{i}-1\bigr{)}+L\epsilon\,.$ (41)
Now, combining all the estimates (28), (29), and (30) gives
$\Phi(A)-\Phi(\tilde{D})<c_{2}\big{(}{\textstyle\sum_{i=1}^{d}}m_{i}-1\big{)}+\epsilon\big{(}\tfrac{3}{2}L+M+M{\textstyle\binom{n}{k}}^{\frac{1}{2}}\big{)}\,.$ (42)
On the other hand, by [DLY24, 2.3] and recalling that $\operatorname{spt}D$ is a unit cube, we have the following lower estimate
$c_{1}\bigl{(}\mathbf{M}(A)-\mathbf{M}(\tilde{D})\bigr{)}\geq c_{1}\bigl{(}\|\gamma_{A}\|_{\mathrm{TV}}-1\bigr{)}\,.$ (43)
Moreover, we observe
${\textstyle\sum_{i=1}^{d}}m_{i}\leq{\textstyle\sum_{i=1}^{d}}m_{i}+{\textstyle\sum_{\lambda\in L}}m_{\lambda}=\|\tilde{\mu}\|_{\mathrm{TV}}\leq\|\tilde{\mu}-\gamma_{A}\|_{\mathrm{TV}}+\|\gamma_{A}\|_{\mathrm{TV}}<\epsilon+\|\gamma_{A}\|_{\mathrm{TV}}\,;$
thus,
$c_{1}\big{(}\|\gamma_{A}\|_{\mathrm{TV}}-1\big{)}=c_{1}\big{(}\|\gamma_{A}\|_{\mathrm{TV}}+\epsilon-\epsilon-1\big{)}>c_{1}\bigl{(}{\textstyle\sum_{i=1}^{d}}m_{i}-1\bigr{)}-c_{1}\epsilon\,.$ (44)
Finally, since $(A,\tilde{D})$ is a polyhedral test pair and $F\in\mathrm{AUE}(\mathcal{P},c_{1})$, combining (42), (43), and (44) we obtain
$c_{2}\bigl{(}{\textstyle\sum_{i=1}^{d}}m_{i}-1\bigr{)}+\epsilon\bigl{(}\tfrac{3}{2}L+M+M{\textstyle\binom{n}{k}}^{\frac{1}{2}}\bigr{)}\\ >\Phi(A)-\Phi(\tilde{D})\geq c_{1}\bigl{(}\mathbf{M}(A)-\mathbf{M}(\tilde{D})\bigr{)}>c_{1}\bigl{(}{\textstyle\sum_{i=1}^{d}}m_{i}-1\bigr{)}-c_{1}\epsilon\,.$ (45)
Recalling the definition of $\epsilon$ we see that this cannot happen. ∎
5.11 Remark.
Note that if $F$ is of class $\mathscr{C}^{1}$, then standard tools (see [Fed69, 3.1.14]) allow to extend $F$ to a $\mathscr{C}^{1}$ function on some compact neighbourhood of $\mathbf{G}_{0}(n,k)$ and then it is clear that $F$ is Lipschitzian so our theorem applies to any integrand of class $\mathscr{C}^{1}$.
Acknowledgements
This research has been supported by National Science Centre Poland grant number 2022/46/E/ST1/00328. The author would like to extend sincere gratitude to Sławomir Kolasiński for his invaluable assistance and insightful discussions throughout this research.